# AS3D: 2D-Assisted Cross-Modal Understanding with Semantic-Spatial Scene Graphs for 3D Visual Grounding

### v2

@@ Abstract
3D visual grounding aims to localize the unique target described by natural languages in 3D scenes. The significant gap between 3D and language modalities makes it a notable challenge to distinguish multiple similar objects through the described spatial relationships. Current methods attempt to achieve cross-modal understanding in complex scenes via a target-centered learning mechanism, ignoring the perception of referred objects. We propose a novel 2D-assisted 3D visual grounding framework that constructs semantic-spatial scene graphs with referred object discrimination for relationship perception. The framework incorporates a dual-branch visual encoder that utilizes 2D pre-trained attributes to guide the multi-modal object encoding. Furthermore, our cross-modal interaction module uses graph attention to facilitate relationship-oriented information fusion. The enhanced object representation and iterative relational learning enable the model to establish effective alignment between 3D vision and referential descriptions. Experimental results on the popular benchmarks demonstrate our superior performance compared to state-of-the-art methods, especially in addressing the challenges of multiple similar distractors.  [^1] The code will be released at https://github.com/onmyoji-xiao/AS3D.
Index Terms:
@@
3D visual grounding, vision-language model, point cloud, scene graph, relationship perception,
I Introduction
@@
Visual grounding plays an important role in 3D scene understanding, applied in various fields including robot navigation [1], autonomous driving [2], and human-computer interaction [3]. Given the textual description and 3D scene, 3D visual grounding aims to localize the unique target corresponding to the description. The main challenge lies in differentiating similar candidate objects within the same scene. It necessitates effective alignment between referential relationships in natural languages and the spatial positions in real scenes.
Figure 1: Previous 3D visual grounding methods can be summarized as predicting the target by the multi-modal fusion of 3D vision and text after respective encoding. Differently, we leverage pre-trained single-view features to guide object encoding and object-level 3D-language alignment, and increase explicit prediction of referred objects to achieve relational learning based on a scene graph.
@@
Recent methods for 3D visual grounding attempt to independently encode the features of 3D point clouds and text into a shared feature space for 3D-language alignment [4]. Since He et al. [5] introduce the Transformer [6] architecture to establish entity-and-relation aware attention, an increasing number of researchers are focusing on utilizing Transformer-based cross-modal fusion to enhance the interaction between visual and textual information [7, 8, 9, 10, 11]. In the latest methods, Xu et al. [12] and Chang et al. [13] respectively incorporate text-related attribute awareness and spatial awareness into the Transformer to achieve more accurate target localization. However, distinguishing targets from intra-class distractors remains challenging due to insufficient relational reasoning across the significant modality gap between 3D vision and language modalities. As shown in Figure 1, existing methods heavily rely on the target probabilities output from the grounding head during training, ignoring the critical influence of the referred objects under inferential conditions. In this paper, we aim to develop a 3D visual grounding framework that promotes multi-modal representations of key objects and relationship perception. This should improve the differentiation of similar objects within complex scenes.
@@
Understanding spatial relationships in 3D scenes can be divided into two primary aspects: (i) identifying objects with attributes and semantic categories corresponding to linguistic targets and referred objects; (ii) inferring the spatial structure between objects that satisfy the described relations. The sparse, noisy, and incomplete characteristics of 3D point clouds lead to the lack of color information and texture details [14], preventing them from fulfilling the object encoding in (i). Although many studies encode 2D views for a more complete representation, a gap remains between 3D vision and languages, which somewhat restricts semantic alignment. Regarding (ii), the scene graph offers a structured representation of the scene and enables reasoning by allowing explicit modeling of object relationships and attributes [15]. However, previous works only focus on neighboring node relationships [16, 17], without gaining significant advantages over other relation-aware structures.
@@
In this paper, we further explore 2D-assisted learning and scene graph construction in the 3D visual grounding task and propose a novel framework, AS3D, to address the challenge of multiple distractors. The primary innovations include a dual-branch 3D encoder that requires only single-view attribute guidance, a semantic-spatial scene graph based on referred objects, and iterative relational learning for multi-modal interaction. Rather than directly concatenating 2D projection features during point cloud encoding, we leverage the pre-trained 2D multi-modal model to improve 3D object-level representation and design a dual-branch network. We utilize interactive contrastive learning among the initial geometric features, class-aware prompts, and CLIP [18] embeddings to improve semantic alignment with limited 3D training data. Subsequently, the referred objects and candidate targets undergo a preliminary screening by aligning the class-aware prompts with object features derived from multi-modal encoding. These objects are then used as nodes to construct a semantic-spatial scene graph. We integrate graph attention layers into the cross-modal interaction module, mainly driven by cross-attention mechanisms. The model updates node information through graph aggregation by iteratively perceiving relevant spatial relationships in the descriptions. The grounding head then outputs the target with the highest probability as the localization result. We evaluate our method on the widely used ReferIt3D [19] and ScanRefer [20] benchmarks, and it outperforms the existing state-of-the-art methods.
@@
  • We propose a 2D-assisted 3D visual grounding framework based on a semantic-spatial scene graph. This framework aims to improve localization in the presence of multiple distractors through more accurate referred object discrimination and awareness of multi-modal relationships.
  • We design a dual-branch 3D visual encoder that employs a pre-trained 2D multi-modal model for feature guidance and assisted supervision. This establishes a robust foundation for the effective construction of scene graphs.
  • Our method is the first work to introduce iterative graph relation perception into vision-language interaction modules, establishing a novel relation-aware cross-modal understanding paradigm.
II Related Work
II-A 3D Visual Grounding
@@ II-A1 Two-stage Localization
The two-stage architecture typically employs dedicated instance segmentation or detection models to generate candidate object proposals, such as VoteNet [21], PointGroup [22], and GroupFree [23]. Existing methods propose distinct 3D visual perception modules and integrate textual features with visual features to identify optimal prediction targets. Recent studies reveal two main challenges: spatial description ambiguity due to variations in viewpoint, and the matching of point clouds and texts while dealing with interference from similar objects. LanguageRefer [8] and VPP-Net [24] adapt viewpoint-dependent descriptions by predicting the orientation of observing viewpoints to adjust the position of point clouds. On the other hand, MVT [9] and ViewRefer [10] encode multiple rotating scene information into visual features simultaneously to implicitly learn viewpoint consistency. For cross-modal feature alignment, FFL-3DOG [17] and TGNN [16] construct a scene graph and measure the relevance of text embeddings and nodes to localize the target. MiKASA [13] presents the multi-key-anchor technique of text-spatial-object fusion encoding to enhance spatial understanding. Although the methods mentioned above achieve remarkable performance, they do not highlight the localization of semantically related referred objects, making relationship learning in complex scenes still challenging.
II-A2 2D-3D Vision Fusion
@@
Sparse geometric representations derived from raw point clouds inherently lack the capacity to capture fine-grained surface textures and semantic attributes. This limitation has prompted existing frameworks to integrate 2D visual information with 3D point cloud features for enhanced 3D localization accuracy. Semantic segmentation-based approaches [20, 25, 26, 27] utilize lightweight architectures like ENet [28] to establish dense 2D-3D correspondence by orthogonally projecting pixel-wise semantic features onto point clouds. 3DJCG [29] and MA2TransVG [12] project the multi-view features to object proposals that are subsequently fused with geometric primitives to form object attribute representations. SAT [7] utilizes a pre-trained Faster-RCNN [30] detector to extract the region of interest features from sampled frames for 2D semantically assisted alignment. LAR [31] synthesizes the 2D images from virtual cameras at five principal viewing directions of object proposals and extracts the image features fused with the 3D features. Our method derives the object attribute from just one view to enhance the 3D point cloud encoding and attain improved multi-modal alignment.
II-B Vision-Language Pretraining
@@
Pre-trained vision-language models learn joint visual and linguistic representations on large-scale data, widely used for downstream tasks related to multi-modal scene understanding [32]. In the image-text domain, CLIP [18] is the most representative pre-trained model for cross-modal semantic alignment of image and text through contrast learning. ALIGN [33] enhances generalization in long-tailed scenarios by scaling up contrastive learning through massive data augmentation. In contrast, ALBEF [34] employs a dual-stream architecture integrated with multi-task learning to strengthen cross-modal alignment through joint representation optimization. Furthermore, researchers also extend visual language pre-training to other application scenarios using contrastive learning as the core foundation [35, 36, 37].
@@
In 3D multi-modal tasks, the scarcity of 3D data and costly annotation pose significant challenges for training on large-scale real-world datasets [38]. Recent approaches leverage pre-trained 2D multi-modal models and conversion relationships between 2D and 3D visual representations to enhance cross-modal perception of 3D vision and languages in 3D-input networks. ULIP [39] uses frozen image and text encoders from CLIP to construct a unified representation of image, text, and 3D point cloud to improve the performance of 3D backbone networks. SemAbs [40] solves the open-world 3D scene understanding problem by combining the semantic abstraction capabilities of 2D visual language models with 3D spatial and geometric reasoning. In this paper, we leverage the pre-trained knowledge of a 2D vision-language model to improve the alignment of point cloud features with text features, while incorporating multi-modal attributes as supplementary information to enhance the 3D visual representation.
@@
Figure 2: The overall architecture of our proposed AS3D. The input scene point clouds are segmented into object point clouds, which are then processed by a dual-branch object encoder. We leverage 2D multi-modal features from the frozen CLIP model to guide additional feature encoding. Object-level feature alignment is calculated exclusively during the training. Instance identifiers are separated from the original textual descriptions and expanded into class-aware semantic prompts. This enables precise semantic correspondences, which then create a relational graph structure that captures dependencies between objects. Cross-modal fusion is achieved through an iterative architecture that integrates cross-attention mechanisms with graph attention layers. The grounding head implemented by the MLP outputs the target with the highest probability.
Figure 3: The network structure of the cross-modal interaction module. The network mainly consists of cross-attention layers, graph attention layers, and geometry-related fusion layers. The content inside the gray box is iterated twice.
III Methods
@@
As shown in Figure 2, our framework includes a 2D-assisted dual-branch 3D object encoder, a semantic-spatial scene graph construction module, and a cross-modal interaction module. The 3D point cloud network performs sampling and encoding for each object proposal in the scene. It utilizes a dual-branch architecture that supports 2D-assisted 3D vision-language alignment and attribute-guided feature extraction. For benchmarks without pre-defined object bounding boxes, we employ the class-agnostic object segmentation model, MaskClustering [41], to attain instance segmentation free from semantic categories. The following scene graph construction improves object relational understanding through semantic matching and the integration of spatial priors. The cross-modal interaction module computes object-text cross-attention and graph attention alternately, ultimately predicting the target with the highest correlation to the description.
III-A Multi-Modal 3D Object Encoding
@@
Existing works mainly encode the point clouds of objects as pure visual features before performing cross-modal interactions. Due to insufficient 3D data and scenes, it is challenging for 3D vision and language information to achieve the same alignment performance as 2D-language learning trained on large-scale data. To address this problem, we utilize the guidance of aligned 2D multi-modal attributes in visual encoding alongside point-level training. Consequently, we propose a dual-branch point cloud network capable of extracting 3D multi-modal features.
III-A1 3D Features with 2D-Assisted Alignment
@@
Our goal is to align object-level 3D vision with language using an open vocabulary. This is accomplished by integrating auxiliary supervision from large-scale 2D multimodal models during the training phase. Specifically, we leverage the pre-trained multi-modal architecture, CLIP [18], to jointly encode visual representations and textual descriptions. The image branch of CLIP extracts 2D visual features from object-bounding boxes projected onto randomly sampled video frames, employing a single randomly selected feature per object for auxiliary training. The text branch, on the other hand, receives class-aware prompts formulated using a template: “The object is [instance name]”, from which textual features are derived. These instance names correspond to the object types found in the training set utterances. Following this, the 3D point cloud features are aligned with textual modalities by optimizing a contrastive loss. This loss aims to preserve both the cross-modal similarity and the inherent semantic correspondence present in CLIP. Taking inspiration from CLIP-KD [44], the visual and textual features from the pre-trained model are further distilled into the 3D vision and text understanding tasks through interactive contrastive learning. The 3D object-text alignment loss $\mathcal{L}_{ot}$ with 2D assistant training is defined as:
@@
III-A2 Attribute-Guided Feature Extraction
to represent the color and texture attributes of object surfaces adequately Sparsely sampled point clouds often lack the ability to represent the color and texture attributes of object surfaces adequately. To address this limitation, many studies extract 2D-view features to achieve a richer representation [20, 12, 46]. However, the process of point-level feature fusion necessitates pixel-level encoding and decoding on the projection of complete views, resulting in substantial computational costs. In contrast, our approach solely extracts object features from the most representative view and integrates them into the geometric encoding of the point cloud network. Distinct from conventional separate encoding schemes, our method seamlessly incorporates 2D attributes as guidance within the intermediate layers of the point cloud network. The cross-modal feature fusion process is formulated as:
@@ III-A2 Attribute-Guided Feature Extraction
Our dual-branch point cloud encoding network ultimately outputs geometrically encoded features $F_{P}$ and 2D attribute-guided multi-modal features $F_{M}$. $F_{P}$ is trained with extra alignment loss and added to $F_{M}$ as object features $F_{O}$. Notably, our framework exclusively processes the single view with the maximal projection coverage for each object, where the projection coordinates are directly derived from the preliminary segmentation outputs. This view selection mechanism, coupled with the direct utilization of pre-segmented spatial mappings, substantially reduces computational overhead by eliminating redundant viewpoint sampling and geometric reconstruction steps.
III-B Semantic-Spatial Scene Graph

### v3

# LSVG: Language-Guided Scene Graphs with 2D-Assisted Multi-Modal Encoding for 3D Visual Grounding

@@ Abstract
3D visual grounding aims to localize the unique target described by natural languages in 3D scenes. The significant gap between 3D and language modalities makes it a notable challenge to distinguish multiple similar objects through the described spatial relationships. Current methods attempt to achieve cross-modal understanding in complex scenes via a target-centered learning mechanism, ignoring the modeling of referred objects. We propose a novel 3D visual grounding framework that constructs language-guided scene graphs with referred object discrimination to improve relational perception. The framework incorporates a dual-branch visual encoder that leverages pre-trained 2D semantics to enhance and supervise the multi-modal 3D encoding. Furthermore, we employ graph attention to promote relationship-oriented information fusion in cross-modal interaction. The learned object representations and scene graph structure enable effective alignment between 3D visual content and textual descriptions. Experimental results on popular benchmarks demonstrate our superior performance compared to state-of-the-art methods, especially in handling the challenges of multiple similar distractors.  [^1] The code will be released at https://github.com/onmyoji-xiao/LSVG.
keywords:
@@
3D visual grounding , scene graph , vision-language model , point cloud , relational perception
journal: Pattern Recognition
[…]
1 Introduction
@@
Visual grounding plays an important role in 3D scene understanding, applied in various fields including robot navigation, autonomous driving, and human-computer interaction [1]. Given the textual description and 3D scene, 3D visual grounding aims to localize the unique target corresponding to the description. The main challenge lies in differentiating similar candidate objects within the same scene [2]. It necessitates effective alignment between referential relationships in natural languages and the spatial positions in real scenes.
Fig. 1: An example of text-guided localization in a scene. In the text, the target is marked in yellow and the referred object is marked in blue. Previous methods (a) construct a fully connected scene graph or use KNN to find objects with spatial neighbors for connection. Instead, our method (b) creates multi-modal representations for all objects and then constructs the scene graph by connecting objects that are identified as referred objects or targets through linguistic cues analysis.
@@
Recent methods for 3D visual grounding attempt to independently encode the features of 3D point clouds and text into a shared feature space for 3D-language alignment [1]. He et al. [3] first introduce the Transformer [4] architecture to establish entity-and-relation aware attention, then an increasing number of researchers are focusing on utilizing Transformer-based fusion to achieve the cross-modal interaction [5, 6, 7, 8, 9]. In the latest methods, Xu et al. [10] and Chang et al. [11] respectively incorporate text-related attribute awareness and spatial awareness into the Transformer to achieve more accurate target localization. However, distinguishing targets from intra-class distractors remains challenging due to insufficient relational reasoning between 3D vision and language modalities. The scene graph provides a structured representation of the scene and allows reasoning by explicitly modeling object relationships and attributes [12]. It plays a pivotal role in visual scene comprehension and has been extensively investigated in existing research [13]. However, prior methods typically construct fully connected graphs [14, 11, 15] or rely solely on KNN (K-Nearest Neighbors) for edge construction [16, 17]. These designs lead to high computational costs due to too many nodes or fail to capture long-range relations. Treating all objects equally can also make it hard to distinguish similar relationships.
@@
Understanding spatial relationships in 3D scenes can be divided into two primary aspects: (i) identifying objects with attributes such as color, shape, scale, and semantic categories corresponding to linguistic targets and referred objects; (ii) inferring the spatial structure between objects that satisfy the described relationships. The sparse, noisy, and incomplete characteristics of 3D point clouds lead to the lack of color information and texture details [18], preventing them from fulfilling the object encoding in (i). Although many studies [19, 20, 21, 22, 23, 24] encode 2D views for a more complete representation, they fail to bridge the semantic gap arising from the modality discrepancy between 3D visual and natural language representations. Instead of existing methods, our approach confines the scope of relational perception to the area between linguistically referred objects, as shown in Fig. 1 (b). 2D views are not only used to enhance point cloud representation but also to supervise 3D-language alignment. We aim to reinforce the crucial relations within the scene graph by creating multi-modal representations during the point cloud encoding process.
@@
In this paper, we further explore scene graph construction in the 3D visual grounding task and propose a novel framework, LSVG, to address the challenge of multiple distractors. The primary innovations include a dual-branch 3D multi-modal encoder that requires only single-view guidance, and a language-guided scene graph on predicted referred objects. Rather than directly concatenating 2D projection features during point cloud encoding, we leverage the pre-trained 2D multi-modal model to enhance and supervise 3D visual encoding. The potential referred objects undergo a preliminary screening by aligning the class-aware prompts with object features derived from multi-modal encoding. The scene graphs with these objects as nodes only perform cross-modal information aggregation and relational reasoning for the relationship clues in the textual descriptions. The target node exhibits strong activation from its contextual connection with the correct referred objects and attains the highest prediction score. We evaluate our method on the widely used ReferIt3D [2] and ScanRefer [19] benchmarks, and it outperforms the existing state-of-the-art methods.
@@
  • We propose a new framework for 3D visual grounding via language-guided scene graphs. The framework aims to improve localization accuracy in complex scenes by enabling more precise identification of referred objects and relation-aware contextual reasoning with graph attention.
  • We design a dual-branch 3D visual encoder that employs a pre-trained 2D multi-modal model for feature enhancement and multi-modal supervision. This establishes a robust foundation for the effective construction of scene graphs.
  • Unlike prior works that rely on fully-connected or KNN-based graph structures, our scene graph edges are based on 3D-language semantic matching. This approach directly connects nodes that are potential referred objects or targets, reducing noise and computational overhead.
2 Related Work
2.1 3D Visual Grounding
@@ 2.1.1 Two-stage Localization
The two-stage architecture typically employs dedicated instance segmentation or detection models to generate potential object proposals, such as VoteNet [25], PointGroup [26], and GroupFree [27]. Existing methods propose distinct 3D visual perception modules and integrate textual features with visual features to identify optimal prediction targets. Recent studies reveal two main challenges: spatial description ambiguity due to variations in viewpoint, and the matching of point clouds and texts while dealing with interference from similar objects. LanguageRefer [6] and VPP-Net [28] adapt viewpoint-dependent descriptions by predicting the orientation of observing viewpoints to adjust the position of point clouds. On the other hand, MVT [7] and ViewRefer [8] encode multiple rotating scene information into visual features simultaneously to implicitly learn viewpoint consistency. Cross-modal feature alignment methods typically employ scene graphs to model text-to-object relevance. FFL-3DOG [14] and TGNN [16] measure text-node similarity through graph structures, while MiKASA [11] introduces multi-key-anchor fusion encoding for object-pair modeling. In contrast, R2G [15] constructs semantic concept-based graphs using heuristic rules. Although the methods mentioned above achieve remarkable performance, they do not highlight the localization of referred objects, making relational learning in complex scenes still challenging.
2.1.2 2D-3D Vision Fusion
@@
Recent advances in 3D visual grounding have demonstrated the benefits of incorporating 2D visual features to complement the sparse geometric information in point clouds. Semantic segmentation-based approaches [19, 20, 23, 21] utilize lightweight architectures like ENet [29] to establish dense 2D-3D correspondence by orthogonally projecting pixel-wise features onto point clouds. 3DJCG [22] and MA2TransVG [10] project the multi-view features to object proposals that are subsequently fused with geometric primitives to form attribute representations. SAT [5] utilizes a pre-trained Faster-RCNN [30] detector to extract the region of interest features from sampled frames for 2D semantically assisted alignment. LAR [31] synthesizes the 2D images from virtual cameras at five principal viewing directions of object proposals and extracts the features. Directly concatenating 2D features with 3D point clouds disrupts the original 2D-language alignment learned by vision-language models. Recent works like CADN [32] highlight the importance of explicit cross-modal consistency in 2D visual grounding. Inspired by this, we extend such alignment to 3D scenes.
Our method derives the object features from just one view to enhance the 3D point cloud encoding and attain improved multi-modal alignment.
2.2 Vision-Language Pretraining
@@
Pre-trained vision-language models learn joint visual and linguistic representations on large-scale data, widely used for downstream tasks related to multi-modal scene understanding [33]. In the image-text domain, CLIP [34] is the most representative pre-trained model for cross-modal semantic alignment of image and text through contrast learning. ALIGN [35] enhances generalization in long-tailed scenarios by scaling up contrastive learning through massive data augmentation. In contrast, ALBEF [36] employs a dual-stream architecture integrated with multi-task learning to strengthen cross-modal alignment through joint representation optimization. Furthermore, researchers also extend visual language pre-training to other application scenarios using contrastive learning as the core foundation [37].
@@
In 3D multi-modal tasks, the scarcity of 3D data and costly annotation pose significant challenges for training on large-scale real-world datasets [38]. Recent approaches leverage pre-trained 2D multi-modal models and conversion relations between 2D and 3D visual representations to enhance cross-modal perception of 3D vision and languages in 3D-input networks. ULIP [39] uses frozen image and text encoders from CLIP to construct a unified representation of image, text, and 3D point cloud to improve the performance of 3D backbone networks. In this paper, we leverage the pre-trained knowledge of a 2D vision-language model to improve the alignment of point cloud features with text features, while incorporating multi-modal features as supplementary information to enhance the 3D visual representation.
@@
Fig. 2: The overall architecture of our proposed LSVG. The input scene point clouds are segmented into object point clouds, which are then processed by a dual-branch object encoder. We leverage 2D multi-modal features from the frozen CLIP model to guide additional feature encoding. Object-level feature alignment is calculated exclusively during training. Instance identifiers are separated from the original textual descriptions and expanded into class-aware semantic prompts. This enables precise 3D-language correspondences, which then create a scene graph that captures dependencies between objects. Cross-modal interaction is achieved through an iterative graph attention architecture. The grounding head implemented by the MLP outputs the target with the highest probability.
3 Methods
@@
As shown in Figure Fig. 2, our framework includes a 2D-assisted dual-branch 3D object encoder, a language-guided scene graph, and a cross-modal interaction module. The 3D point cloud network performs sampling and encoding for each object proposal in the scene. It utilizes a dual-branch architecture that supports 2D-supervised 3D-language alignment and feature enhancement. For benchmarks without pre-defined object bounding boxes, we employ the class-agnostic object segmentation model, MaskClustering [40], to attain instance segmentation free from semantic categories. The scene graph is constructed with the semantically relevant candidates as nodes, which are extracted through 3D-language matching. The cross-modal interaction module alternately performs object-text cross-attention and graph attention, ultimately predicting the target with the highest correlation to the description.
3.1 Multi-Modal 3D Object Encoding
@@
Existing works mainly encode the point clouds of objects as pure visual features before performing cross-modal interactions. Due to insufficient 3D data and scenes, it is challenging for 3D vision and language information to achieve the same alignment performance as 2D-language learning trained on large-scale data. To address this problem, we utilize the guidance of aligned 2D multi-modal features in visual encoding alongside point-level training. Consequently, we propose a dual-branch point cloud network capable of extracting 3D multi-modal features.
3.1.1 2D-Supervised 3D-Language Alignment
@@
Our goal is to align object-level 3D vision with language using an open vocabulary. This is accomplished by integrating auxiliary supervision from large-scale 2D multi-modal models during training. Specifically, we leverage the pre-trained CLIP [34] to jointly encode visual representations and textual descriptions. The image branch of CLIP extracts 2D visual features from object-bounding boxes projected onto randomly sampled video frames, employing a single randomly selected feature per object for auxiliary training. The text branch, on the other hand, receives class-aware prompts formulated using a template: “The object is [instance name]”, from which textual features are derived. These instance names correspond to the object types found in the training set utterances. Following this, the 3D point cloud features are aligned with textual modalities by optimizing a contrastive loss. This loss aims to preserve both the cross-modal similarity and the inherent semantic correspondence present in CLIP. Taking inspiration from CLIP-KD [42], the visual and textual features from the pre-trained model are further distilled into the 3D vision and text understanding tasks through interactive contrastive learning. The 3D object-text alignment loss $\mathcal{L}_{ot}$ with 2D assistant training is defined as:
@@
3.1.2 2D-Enhanced Feature Extraction
Sparsely sampled point clouds often lack the ability to represent the color and texture attributes of object surfaces adequately. To address this limitation, many studies extract 2D-view features to achieve a richer representation [19, 10, 44]. However, the process of point-level feature fusion necessitates pixel-level encoding and decoding on the projection of complete views, resulting in substantial computational costs. In contrast, our approach solely extracts object features from the most representative view and integrates them into the geometric encoding of the point cloud network. Distinct from conventional separate encoding schemes, our method seamlessly incorporates 2D features as enhancement within the intermediate layers of the point cloud network. The cross-modal feature fusion process is formulated as:
@@
Our dual-branch point cloud encoding network ultimately outputs geometrically encoded point cloud features $F_{P}$ and 2D-enhanced object features $F_{M}$. $F_{P}$ is trained with extra alignment loss and added to $F_{M}$ as object features $F_{O}$. Notably, our framework exclusively processes the single view with the maximal projection coverage for each object, where the projection coordinates are directly derived from the preliminary segmentation outputs. This view selection mechanism, coupled with the direct utilization of pre-segmented spatial mappings, substantially reduces computational overhead by eliminating redundant viewpoint sampling and geometric reconstruction steps.
Fig. 3: The network structure of the cross-modal interaction module. The network mainly consists of cross-attention layers, graph attention layers, and geometry-related fusion layers. The content inside the gray box is iterated twice.
3.2 Language-Guided Scene Graph